# Gradient correlation is needed to accelerate SGD with momentum

### v1

@@ A.2 Neural networks
As a complement, we display on Figure 6 a slightly different view of Figure 2. On the left part of Figure 6(a)-Figure 6(b), we display the typical behaviour of one single run of optimization algorithms, namely without averaging for several initialisations. As we can expect, we observe a higher variability, although the general behaviour remains similar. On the right part of Figure 6(a)-Figure 6(b), we displayed the evolution of RACOGA values along the iterations, instead of the histogram distribution of Figure 2, which does not keep any temporal information. For correlated data, we observe that RACOGA values decrease when iterations converge. This phenomenon can be interpreted as the convergence of the iterations to a minimum with low curvature which is related to a low RACOGA value (see Appendix H for more details).
@@
Figure 7: Illustration of the convergence speed of GD (Algorithm 6), SGD (Algorithm 1, batch size $64$), NAG (Algorithm 8) and SNAG (Algorithm 3, batch size $64$), ADAM (Algorithm 5) and RMSprop (Algorithm 4) averaged over $10$ different initializations, together with an histogram distribution of RACOGA values taken along the optimization path, averaged over $10$ different initialisations. On the left, we use a MLP to classify fewly correlated data sampled from an isotropic law. On the right we use a CNN to classify CIFAR10 images. Note that if ADAM and RMSprop are both better than other algorithms for the SPHERE experiment, they are not faster than SNAG in the CIFAR10 experiment.
@@
Using the Python library time, the computation time needed to choose the best parameters for our two models (CNN for CIFAR10 and MLP for SPHERE) have been saved. Moreover, the computational time needed to generate Figures 123 and 7 is added to our computational budget. We saved the computation time needed to train the models and to compute the RACOGA values, $10$ times per algorithms (due to the $10$ initializations) and for both networks. Note that for the stochastic algorithms, RACOGA was computed only every $100$ iterations, because of the heavy computation time it demands. The experiments for the linear regression problem take less than one minute of computation. Experiments with SNAG, NAG, SGD and GD took approximately $4$ hours. Additional experiments with ADAM and RMSprop took approximately $8$ hours. Note that ADAM requires to tune 3 hyperparameters with the grid search, making this step significantly longer. The total computation time needed for all these experiments is approximately $12$ hours of computation on a GPU NVIDIA A100 Tensor Core.
@@ B.2 The different forms of Nesterov algorithm
Appendix C Related results
@@ C.1 Gradient correlation conditions
However, RACOGA appears naturally in our study of the SGC, and has the advantage to make explicitly appear the correlation between gradients.
@@ E.1 Proofs of Theorem 1 and Theorem 2
We can deduce Theorem 1-Theorem 2 by respectively adapting the proof from Garrigos & Gower (2023) and Sebbouh et al. (2021). We can not directly apply their original results because their setting is slightly different: in problem (FS) their convexity assumption holds for each $f_{i}$, while in our case it solely holds over the whole sum. In both expectation and almost sure cases, the core Lemma is to bound $\mathbb{E}\left[\lVert\tilde{\nabla}_{K}(x)\rVert^{2}\right]$, which bounds the variance of the estimator. It is Lemma 6.7 in Garrigos & Gower (2023)), and Lemma 1.3 in Sebbouh et al. (2021). These two lemmas do not hold in our setting. We will use instead the following result.
@@ E.2 Convergence of SGD with strong growth condition
Theorem 1 states a convergence result for SGD (Algorithm 1) without making use of the strong growth condition. It is however possible, and we do so in Theorem 8.
@@
Appendix F Convergence of SNAG with strong growth condition
In Sections F.1 and F.2, we provide for completeness a proof of Theorem 3, that is a similar result to the one from Vaswani et al. (2019). Our proof is a slightly simpler formulation of the algorithm. In Sections F.3 and F.4, we extend these results proving new almost sure convergences (Theorem 4), that are asymptotically better that the results in expectation.

### v2

@@ A.2 Neural networks
As a complement, we display on Figure 7 a slightly different view of Figure 2. On the left part of Figure 7(a)-Figure 7(b), we display the typical behaviour of one single run of optimization algorithms, namely without averaging for several initialisations. As we can expect, we observe a higher variability, although the general behaviour remains similar. On the right part of Figure 7(a)-Figure 7(b), we displayed the evolution of RACOGA values along the iterations, instead of the histogram distribution of Figure 2, which does not keep any temporal information. For correlated data, we observe that RACOGA values decrease when iterations converge. This phenomenon can be interpreted as the convergence of the iterations to a minimum with low curvature which is related to a low RACOGA value (see Appendix I for more details).
@@
Figure 8: Illustration of the convergence speed of GD (Algorithm 6), SGD (Algorithm 1, batch size $64$), NAG (Algorithm 3, full batch) and SNAG (Algorithm 3, batch size $64$), ADAM (Algorithm 5) and RMSprop (Algorithm 4) averaged over $10$ different initializations, together with an histogram distribution of RACOGA values taken along the optimization path, averaged over $10$ different initialisations. On the left, we use a MLP to classify fewly correlated data sampled from an isotropic law. On the right we use a CNN to classify CIFAR10 images. Note that if ADAM and RMSprop are both better than other algorithms for the SPHERE experiment, they are not faster than SNAG in the CIFAR10 experiment.
@@
Using the Python library time, the computation time needed to choose the best parameters for our two models (CNN for CIFAR10 and MLP for SPHERE) have been saved. Moreover, the computational time needed to generate Figures 123 and 8 is added to our computational budget. We saved the computation time needed to train the models and to compute the RACOGA values, $10$ times per algorithms (due to the $10$ initializations) and for both networks. Note that for the stochastic algorithms, RACOGA was computed only every $100$ iterations, because of the heavy computation time it demands. The experiments for the linear regression problem take less than one minute of computation. Experiments with SNAG, NAG, SGD and GD took approximately $4$ hours. Additional experiments with ADAM and RMSprop took approximately $8$ hours. Note that ADAM requires to tune 3 hyperparameters with the grid search, making this step significantly longer. The total computation time needed for all these experiments is approximately $12$ hours of computation on a GPU NVIDIA A100 Tensor Core.
Additional datasets results
On Figure 9, we present the convergence curves and RACOGA values for SGD (Algorithm 1) and SNAG (Algorithm 3) on various datasets, including hand-written numbers (MNIST), pictures of clothes (FashionMNIST Xiao et al. (2017)) and Kuzushiji characters (KMNIST Clanuwat et al. (2018)) and hand-written characters (EMNIST Cohen et al. (2017)). The used learning rate for SGD is $s=0.1$ for MNIST and EMNIST and $s=0.25$ for FashionMNIST and KMNIST. The learning rate of SNAG is $s=0.05$ for MNIST and EMNIST and $s=0.1$ for Fashion MNIST and KMNIST. The momentum of SNAG (Algorithm 3) is $p=0.8$ for MNIST and FashionMNIST, and $p=0.9$ for KMNIST and EMNIST. These quantities have been chosen after a grid search to find the parameters that maximize the test accuracy of the learned model. There were $3$ epochs for the training.
Note that the RACOGA values are large for all these training paths. Moreover, in all these scenarios, we observe that the convergence of SNAG is faster than the convergence of SGD, although it is less clear for KMNIST. These additional experimental validations support our theoretical results.
(a) MNIST
(b) FashionMNIST
(c) KMNIST
(d) EMNIST
Figure 9: Illustration of the convergence speed of SGD (Algorithm 1 and SNAG (Algorithm 3, batch size $64$), averaged over $10$ different initializations, together with an histogram distribution of RACOGA values taken along one optimization path. We use a CNN described on Figure 6(b), only changing the input layer with the dimension of images $1\times 28\times 28$. Note that SNAG is faster than SGD, although it is less clear for KMNIST, and the RACOGA values are large for all these datasets.
A.3 Logistic regression
On Figure 10, we present the convergence curves and RACOGA values for SGD (Algorithm 1) and SNAG (Algorithm 3) on images extracted from CIFAR10. Images of CIFAR10 have $10$ classes and are of dimension $3,072$, so the logistic regression has $30,730$ parameters. In order to observe the convergence curves on under-parametrized regime, we run the optimization on all the CIFAR10 dataset ($60,000$ images). To analyse the over-parametrized regime, we run the optimization on a subset of $10,000$ images from CIFAR10. The learning rate of SNAG is $s=0.2$ and the momentum $p=0.7$. The learning rate for SGD is $s=1.0$. These parameters have been chosen after a grid search to find the parameters that maximize the test accuracy of the learned model. There were $5$ epochs for the training.
Note that for logistic regression the RACOGA values are positive but small (compared to the deep learning experiments). Therefore, we do not observe any acceleration of SNAG over SGD.
(a) Under-parametrized ($60,000$ data)
(b) Over-parametrized ($10,000$ data)
Figure 10: Illustration of the convergence speed of SGD (Algorithm 1 and SNAG (Algorithm 3, batch size $64$), averaged over $10$ different initializations, together with an histogram distribution of RACOGA values taken along one optimization path. We run a logistic regression ($30,730$ parameters) on images extracted from CIFAR10. Note that SNAG and SGD have the same speed, which may be caused by the non convexity of the problem, or by the fact that the RACOGA values might be not large enough to observe an acceleration of SNAG.
@@ B.2 The different forms of Nesterov algorithm
Appendix C Related works
@@ C.1 Gradient correlation conditions
However, RACOGA appears naturally in our study of the SGC, and is a direct measure of the correlation between gradients.
@@ C.2 Linear regression acceleration results
C.3 SGC related convergence results
The linear convergence of SGD under a variation of SGC, for smooth convex and strongly convex functions, has been addressed in (Schmidt & Roux, 2013). Later, convergence of SNAG type algorithms under strong growth condition, for functions that are convex or strongly convex, also attracted interest. In this section, we discuss how these works relate to Theorems 3-4.
Vaswani et al. (2019)
As mentioned in Section 3, the bounds of Theorem 3 were already achieved by (Vaswani et al., 2019). It was achieved with a quite unusual formulation of the SNAG algorithm, which is equivalent, to the following one
Algorithm 10 (Vaswani et al., 2019))
1:  input: $x_{0}=z_{0}\in{\mathbb{R}}^{d}$, $s>0$, $\beta\in[0,1]$, $(\alpha_{n})_{n\in\mathbf{N}}\in[0,1]^{\mathbf{N}}$, $(\eta_{n})_{n\in\mathbf{N}}\in\mathbf{R}_{+}^{\mathbf{N}}$, $g_{n}(.)$ stochastic approximation of $\nabla f(.)$ at iteration $n$.
2:  for $n=0,1,\dots,n_{it}-1$ do
3:     $y_{n}=\alpha_{n}x_{n}+(1-\alpha_{n})z_{n}$
4:     $x_{n+1}=y_{n}-sg_{n}(y_{n})$
5:     $z_{n+1}=\beta z_{n}+(1-\beta)y_{n}-\gamma_{n}sg_{n}(y_{n})$
6:  end for
7:  output: $x_{n_{it}}$
For $f$ $L$-smooth, and such that SGC is verified with constant $\rho$, the bounds of Theorem 3 are verified considering SNAG (Vaswani et al., 2019) with the following parameters:
$\bullet$ If $f$ $\mu$-strongly convex:
$\gamma_{n}=\frac{1}{\sqrt{\mu s\rho}},\quad\beta_{n}=1-\sqrt{\frac{\mu s}{\rho}},\quad b_{n+1}=\frac{\sqrt{\mu}}{\left(1-\sqrt{\frac{\mu s}{\rho}}\right)^{(n+1)/2}},$
$a_{n+1}=\frac{1}{\left(1-\sqrt{\frac{\mu s}{\rho}}\right)^{(n+1)/2}},\quad\alpha_{n}=\frac{\gamma_{n}\beta_{n}b_{n+1}^{2}s}{\gamma_{n}\beta_{n}b_{n+1}^{n}s+a_{n}^{2}},\quad s=\frac{1}{\rho L}.$
$\bullet$ If $f$ convex:
$\gamma_{n}=\frac{\frac{1}{\rho}+\sqrt{\frac{1}{\rho^{2}}+4\gamma_{n-1}^{2}}}{2},\quad a_{n+1}=\gamma_{n}\sqrt{s\rho},$
$\alpha_{n}=\frac{\gamma_{n}s}{\gamma_{n}s+a_{n}^{2}},\quad s=\frac{1}{\rho L}.$
In the strongly convex case, intermediate sequences of parameters $(a_{n})$ and $(b_{n})$ appear. In the convex case, the sequence $(\gamma_{n})$ is defined with a recursive formula. Our different proof of Theorem 3 does not make these features appear.
Gupta et al. (2023)
Another line of research leads to a similar expectation result to ours, using AGNES (Algorithm 11), and so with a different proof. The authors of Gupta et al. (2023) also get almost sure convergence, nevertheless without convergence rates contrarily to our Theorem 4.
Algorithm 11 Accelerated Gradient descent with Noisy EStimators (AGNES), Gupta et al. (2023)
1:  input:$f$ (objective/loss function), $x_{0}$ (initial point), $\alpha=10^{-3}$ (learning rate), $\eta=10^{-2}$ (correction step size), $\rho=0.99$ (momentum), $N$ (number of iterations)
2:  $v_{0}\leftarrow 0$
3:  for $n=0,1,\dots,N$ do
4:     $g_{n}\leftarrow\nabla_{x}f(x_{n})$  (gradient estimator)
5:     $v_{n+1}\leftarrow\rho(v_{n}-g_{n})$
6:     $x_{n+1}\leftarrow x_{n}+\alpha v_{n+1}-\eta g_{n}$
7:  end for
8:  $g_{N}\leftarrow\nabla_{x}f(x_{N})$
9:  $x_{N}\leftarrow x_{N}-\eta g_{N}$
10:  output: $x_{N}$
@@ E.1 Proofs of Theorem 1 and Theorem 2
We can deduce Theorem 1-Theorem 2 by respectively adapting the proof from Garrigos & Gower (2023) and Sebbouh et al. (2021). We can not directly apply their original results because their setting is slightly different: in problem (FS) their convexity assumption holds for each $f_{i}$, while in our case it solely holds over the whole sum. In both expectation and almost sure cases, the core Lemma is to bound $\mathbb{E}\left[\lVert\tilde{\nabla}_{K}(x)\rVert^{2}\right]$, which bounds the variance of the estimator. It is Lemma 6.7 in Garrigos & Gower (2023), and Lemma 1.3 in Sebbouh et al. (2021). These two lemmas do not hold in our setting. We will use instead the following result.
@@ E.2 Convergence of SGD with strong growth condition
Theorem 1 states a convergence result for SGD (Algorithm 1) without making use of the strong growth condition. It is however possible, as done in Gupta et al. (2023). In Theorem 8 we give a very similar result.
@@
Appendix F Convergence of SNAG without strong growth condition
In this Section, we derive a finite time convergence result in expectation for SNAG (Algorithm 2) without assuming SGC. In this case, as for Theorems 1-2, the bound over the noise is derived from the geometrical properties of the functions.
Lemma 3.
Assume $f$ is such that assumptions (1)-(2) hold.
If $f$ is $\mu$-strongly convex, we have
$\mathbb{E}\left[\lVert\nabla f(x)-\tilde{\nabla}_{K}(x)\rVert^{2}\right]\leq 2(L_{(K)}-\mu)(f(x)-f^{\ast}).$ (66)
If $f$ is convex, we have
$\mathbb{E}\left[\lVert\nabla f(x)-\tilde{\nabla}_{K}(x)\rVert^{2}\right]\leq 2L_{(K)}(f(x)-f^{\ast}).$ (67)
We prove Lemma 3 in Appendix F.3. We note $L$ the smoothness constant of $f$. Under assumption (2), we have $L\leq\frac{1}{N}\sum_{i=1}^{N}L_{i}\leq L_{(K)}$.
Theorem 9.
Under Assumptions 1 and 2, SNAG (Algorithm 2) with batch size $K$ guarantees to reach an $\varepsilon$-precision (2) at the following iterations:
$\bullet$ If $f$ is convex, $s_{n}=\frac{1}{2L_{(K)}}\frac{1}{n+1}$, $\alpha_{n}=\frac{n}{n+2}$, $\eta_{n}=\frac{1}{4}\frac{1}{L_{(K)}}$ , $\beta=1$,
$n\geq\frac{4L_{(K)}}{\varepsilon}\lVert x_{0}-x^{\ast}\rVert^{2}.$ (68)
$\bullet$ If $f$ is $\mu$-strongly convex, $s=\frac{1}{16}\frac{\mu}{(L_{(K)}-\mu)^{2}},~{}\beta=1-\frac{1}{8}\frac{\mu}{(L_{(K)}-\mu)},~{}\eta=\frac{1}{4}\frac{1}{(L_{(K)}-\mu)},~{}\alpha=\frac{1}{1+\frac{1}{4}\frac{\mu}{(L_{(K)}-\mu)}}$,
$n\geq 8\frac{L_{(K)}-\mu}{\mu}\log\left(\frac{2(f(x_{0})-f^{\ast})}{\varepsilon}\right).$ (69)